# Fidelity Isn’t Accuracy: When Linearly Decodable Functions Fail to Match the Ground Truth

### v1

@@ 5.1.2 Medical Insurance Cost
Yet despite the neural network’s superior performance, the linear surrogate fails to retain that advantage: it achieves a lower $R^{2}=0.6659$ against the true target, even while attaining a high linearity score $\lambda(f)=0.9186$. This discrepancy reveals a key insight: the linearly decodable portion of the network’s behavior is not necessarily what makes it accurate. The surrogate mimics the network closely, but what it captures may not be what matters.
@@ 5.1.3 Concrete Compressive Strength
In this dataset, the baseline linear regression model achieves an $R^{2}$ score of 0.5944, reflecting a moderate ability to capture the structure in the data. The neural network substantially improves upon this with $R^{2}=0.8588$, indicating that it successfully models complex nonlinear relationships among the input features.
@@
Figure 4 visualizes the predictions from all three models. The linear surrogate trained to mimic the neural network achieves an $R^{2}$ of 0.5821 against the true target—close to the baseline’s 0.5944—but significantly lower than the network’s 0.8588. Nonetheless, the surrogate yields a linearity score of $\lambda(f)=0.6659$, indicating that it captures over 60% of the variance in the network’s output.
@@ 5.1.4 California Housing
On this dataset, the baseline linear regression model achieves an $R^{2}$ score of 0.5758, indicating moderate linear structure in the input features. The neural network improves on this with $R^{2}=0.7908$, demonstrating that it captures additional nonlinear patterns in the data. The linear surrogate trained to mimic the neural network achieves $R^{2}=0.5658$ against the true target—slightly worse than the baseline—but exhibits relatively high fidelity to the network, with a linearity score of $\lambda(f)=0.6968$.
@@
As in the Concrete dataset, this result reinforces our central claim: a high $\lambda(f)$ reflects how well a neural network’s behavior can be mimicked, but it does not guarantee that the mimic will retain predictive strength. The predictive signal may be concentrated in the nonlinear residuals that a linear surrogate cannot express.
@@ 5.2 Discussion
Our experiments demonstrate that the linearity score $\lambda(f)$ serves as a reliable measure of how linearly decodable a neural network’s output function is. In settings where the learned function is highly nonlinear, such as our synthetic dataset, $\lambda(f)$ drops to zero, confirming that no meaningful linear approximation exists. In contrast, when $\lambda(f)$ is moderate or high—as in the Concrete and California Housing datasets—the surrogate tracks the network’s output with reasonable fidelity, but fails to retain its predictive advantage.
Importantly, a high linearity score indicates fidelity to the neural network, but not necessarily to the ground truth. For example, in the case of the Concrete dataset, approximately 36% of the network’s output variance is not captured by the linear surrogate. However, that residual variance appears to contain nearly all of the model’s predictive improvement over the baseline. The results for the California Housing dataset share this property. These findings suggest that the portion of the neural network’s function that escapes linear approximation—though smaller in variance—may disproportionally contribute to the model’s predictive accuracy.
This distinction highlights an important nuance: a linear surrogate with high $\lambda(f)$ offers insight into the network’s behavior, but may not reproduce its predictive power if that power depends on nonlinear components. In some high-$\lambda$ regimes, the surrogate may still track the network’s output well, but its usefulness as a predictive substitute depends on how much of the model’s performance derives from linearly decodable structure.
While $\lambda(f)$ serves as a useful diagnostic of linear decodability, it is not a comprehensive interpretability method. It offers no insight into a network’s internal mechanisms—such as feature salience, representational geometry, or pathway analysis—nor does it produce explanations in terms of input features, like importance scores or causal attributions. Furthermore, our approach is currently limited to regression tasks; extending it to classification would require adapting the framework to logit or softmax outputs. These limitations suggest that $\lambda(f)$ should be viewed not as a general-purpose interpretability tool, but as a focused, principled diagnostic—particularly well-suited for assessing whether a network’s behavior exhibits linear structure at the output level.
@@ 6 Conclusion
Neural networks are often—and justifiably—described as black boxes, posing challenges in high-stakes domains like healthcare and finance. In this work, we proposed a way to measure just how illuminated or comprehensible a network is by quantifying its linear decodability. The metric, $\lambda(f)$, offers a simple and interpretable means of assessing whether a neural network’s output function is well-approximated by a linear model. It is straightforward to compute and opens the door to further analysis of learned functions in terms of their structural simplicity.
@@
Several promising directions emerge from this work. A natural extension would be to explore how an analogous metric might be applied to classification tasks—such as by projecting logits or softmax outputs—to assess interpretability in non-regression settings. Unlike continuous-valued regression outputs, classification involves discrete decision boundaries, which complicates the notion of linear decodability. However, structure in logit space may still reveal interpretable trends—such as class separability or margin alignment—that a linear surrogate could capture.
@@
Another compelling direction is to move beyond quantifying linear decodability and begin characterizing it. Since $\lambda(f)$ captures the proportion of output variance explained by a linear model, a natural next step is to ask: what components of the network’s behavior are responsible for this? Investigating which input regions, features, or latent representations contribute most to the linearly decodable structure could deepen our understanding of how and where simple functional patterns emerge within otherwise complex models. A similar investigation into the components of the network’s behavior that are not linearly representable is equally—if not more—promising.
@@
Across both synthetic and real-world datasets, our experiments show that $\lambda(f)$ reliably captures how linearly decodable a network’s output function is. However—and this is the central finding of this work—high fidelity to the network’s output does not imply high accuracy on the original prediction task. In several cases, linear surrogates with high $\lambda(f)$ scores closely mimic the neural network’s outputs, yet underperform even a baseline linear model when evaluated against the ground truth. This mismatch reveals that the portions of a network’s behavior that are not linearly approximable—though smaller in variance—often carry the critical predictive signal. As a result, $\lambda(f)$ should be interpreted as a measure of structural alignment with the model, not necessarily with the data-generating process.
@@
By providing a concrete measure of when neural network behavior can be meaningfully approximated by simpler linear models, $\lambda(f)$ represents a step toward more interpretable machine learning systems—systems that acknowledge both the power and the opacity of deep networks while offering a principled diagnostic for understanding their behavior.
Acknowledgments
@@
I thank Professor Jorge Silveyra for a valuable discussion that helped spark this project. I am also grateful for the support of the Lafayette College Computer Science Department.

### v3

@@ 5.1.2 Medical Insurance Cost
Despite the neural network’s superior performance, the linear surrogate fails to retain that advantage: it achieves a lower $R^{2}=0.6659$ against the true target, even while attaining a high linearity score $\lambda(f)=0.9186$. This discrepancy reveals a key insight: the linearly decodable portion of the network’s behavior is not necessarily what makes it accurate. The surrogate mimics the network closely, but the features it captures may not be those that drive predictive performance.
@@ 5.1.3 Concrete Compressive Strength
On the Concrete Compressive Strength dataset, the baseline linear regression model achieves an $R^{2}$ score of 0.5944, reflecting a moderate ability to capture the structure in the data. The neural network substantially improves upon this with $R^{2}=0.8588$, indicating that it successfully models complex nonlinear relationships among the input space.
@@
Figure 4 visualizes the predictions from all three models. The linear surrogate trained to mimic the neural network achieves $R^{2}=0.5821$ against the true target—close to the baseline’s 0.5944—but significantly lower than the network’s 0.8588. Nonetheless, the surrogate yields a linearity score of $\lambda(f)=0.6659$, indicating that it captures over 60% of the variance in the network’s output.
@@ 5.1.4 California Housing
With regard to the California Housing dataset, the baseline linear regression model achieves an $R^{2}$ score of 0.5758, indicating moderate linear structure in the input features. The neural network improves on this with $R^{2}=0.7908$, demonstrating that it captures additional nonlinear patterns in the data. The linear surrogate trained to mimic the neural network achieves $R^{2}=0.5658$ against the true target—slightly worse than the baseline—but exhibits relatively high fidelity to the network, with a linearity score of $\lambda(f)=0.6968$.
@@
As in the Concrete dataset, this result reinforces the key claim that a high $\lambda(f)$ reflects how well a neural network’s behavior can be mimicked, but it does not guarantee that the mimic will retain predictive strength. The predictive signal may be concentrated in the nonlinear residuals that a linear surrogate cannot express.
@@ 5.2 Discussion
The described experiments demonstrate that the linearity score $\lambda(f)$ serves as a reliable measure of the linearly decodablity of a neural network’s output function. In settings where the learned function is highly nonlinear, such as the synthetic dataset, $\lambda(f)$ drops to zero, confirming that no meaningful linear approximation exists. In contrast, when $\lambda(f)$ is moderate or high—as in the Concrete and California Housing datasets—the surrogate tracks the network’s output with reasonable fidelity, but fails to retain its predictive advantage.
$\lambda(f)$ serves as a useful diagnostic of linear decodability, but it is not a comprehensive interpretability method. It offers no insight into a network’s internal mechanisms—such as feature salience, representational geometry, or pathway analysis—nor does it produce explanations in terms of input features, like importance scores or causal attributions. Furthermore, this approach is currently limited to regression tasks; extending it to classification would require adapting the framework to logit or softmax outputs. These limitations make clear that $\lambda(f)$ is not a general-purpose interpretability tool, but rather a focused, principled diagnostic—particularly well-suited for assessing whether a regression network’s behavior exhibits linear structure at the output level.
These results reveal an important distinction: a high linearity score indicates fidelity to the neural network, but not necessarily to the ground truth. For example, in the case of the Concrete dataset, approximately 36% of the network’s output variance is not captured by the linear surrogate. However, that residual appears to contain nearly all of the model’s predictive improvement over the baseline. The results for the California Housing dataset share this property. These findings suggest that the portion of the neural network’s function that escapes linear approximation—though smaller in variance—may disproportionately contribute to the model’s predictive accuracy.
This distinction highlights an important nuance: a linear surrogate with high $\lambda(f)$ offers insight into the network’s behavior, but may not reproduce its predictive power if that power depends on nonlinear components. In some high $\lambda(f)$ regimes, the surrogate may still track the network’s output well, but its usefulness as a predictive substitute depends on how much of the model’s performance derives from linearly decodable structure.
@@ 6 Conclusion
Neural networks are often described as black boxes, posing challenges in high-stakes domains like healthcare and finance. In this work, a metric was proposed to assess the degree to which a network is transparent and comprehensible by quantifying its linear decodability. The metric, $\lambda(f)$, offers a simple and interpretable means of assessing whether a neural network’s output function is well-approximated by a linear model. It is straightforward to compute and facilitates further analysis of learned functions in terms of their structural simplicity.
@@
Several promising directions emerge from this work. A natural extension is the exploration of how an analogous metric might be applied to classification tasks—such as by projecting logits or softmax outputs—to assess interpretability in non-regression settings. Unlike continuous-valued regression outputs, classification involves discrete decision boundaries, which complicates the notion of linear decodability. However, structure in logit space may still reveal interpretable trends—such as class separability or margin alignment—that a linear surrogate could capture.
@@
Another noteworthy direction for future research is to move beyond quantifying linear decodability and begin characterizing it. Since $\lambda(f)$ captures the proportion of output variance explained by a linear model, a natural next step is to determine what components of a network’s behavior are linearly decodable. Investigating which input regions, features, or latent representations contribute most to the linearly decodable structure could deepen the understanding of how and where simple functional patterns emerge within otherwise complex models. A similar investigation into the components of a network’s behavior that are not linearly representable may be equally, if not more, promising.
A final direction is to explore the extent to which these findings generalize. Investigations could examine probes and surrogate models more complex than linear regression, such as kernel methods, tree ensembles, or shallow neural approximators, to test whether the fidelity-accuracy distinction persists under richer function classes. Another important question is whether the same effect holds across different domains, architectures, and training regimes. Establishing these boundaries wold help clarify when fidelity to a network’s behavior meaningfully aligns with predictive accuracy—and when it does not—thereby delineating the true scope of the fidelity-accuracy distinction.
@@
Across both synthetic and real-world datasets, experiments show that $\lambda(f)$ reliably captures how linearly decodable a network’s output function is. However, high fidelity to the network’s output does not imply high accuracy on the original prediction task. In several cases, linear surrogates with high $\lambda(f)$ scores closely mimic the neural network’s outputs, yet underperform even a baseline linear model when evaluated against the ground truth. This mismatch reveals that the portions of a network’s behavior that are not linearly approximable—though smaller in variance—often carry the critical predictive signal. As a result, $\lambda(f)$ should be interpreted as a measure of structural alignment with the model, not necessarily with the data-generating process.
Probes remain a central tool in mechanistic interpretability ferrandoPrimerInnerWorkings2024 , particularly for analyzing representations and circuit-level behavior. However, these findings caution against interpreting probe fidelity as definitive evidence of meaningful alignment with the task or ground truth. In contexts where surrogate interpretability is used as a proxy for scientific insight, faithfulness to the model alone is not enough.
@@
By providing a concrete measure of when neural network behavior can be meaningfully approximated by simpler linear models, $\lambda(f)$ represents a step toward more interpretable machine learning systems—systems that balance the predictive capabilities of deep networks with their interpretability constraints while offering a principled diagnostic for understanding their behavior. As interpretability continues to mature, distinguishing between fidelity and accuracy will be essential for understanding what models truly know.
Acknowledgments and Disclosure of Funding
@@
I thank Professor Jorge Silveyra for a valuable discussion that helped spark this project. I am also grateful for the support of the Lafayette College Department of Computer Science.